An empty Styrofoam cup has a mass of 8.655g. When the cup is completely fill with water (density =0.9971g/ml), the mass is 21.312g. When the cup is emptied of water and then filled with an unknown liquid, the new mass is 29.293g. What is the density of the unknown liquid to the proper number of significant figures? Show your work.
Answer #1

ANSWER:

~Mass of empty styrofoam cup: 8.655 grams

​​​​​​~Mass of completely filled styrofoam cup with water: 21.312 grams

~Mass of water in cup = 21.312 - 8.655 = 12.657

~Now we'll calculate volume of water in cup:

Thus, 12.695 mL of any liquid will be required to completely fill given styrofoam cup.

~ Total mass after filling styrofoam cup with another liquid = 29.293 grams

~ Mass of styrofoam cup = 8.655 grams

~ Mass of unknown liquid = 29.293 - 8.655 = 20.638 grams

~ Volume = 12.695 mL (As volume that can be accommodated in given styrofoam cup for any liquid will be same to completely fill it)

Thus,